# FPT+: A Parameter and Memory Efficient Transfer Learning Method for High-resolution Medical Image Classification

### v1

@@ 3.1 Vision Transformer
Within each self-attention layer, individual tokens in the input sequence are mapped to Q (query), K (key), and V (value) representations for computations in the attention module (Attn) (Vaswani et al., 2017) as follows:
@@
3.2 Side Tuning
Unlike popular PETL methods which introduce supplementary modules into transformer blocks, our approach, drawing inspiration from (Zhang et al., 2020; Sung et al., 2022), adopts a lightweight and separate side network. The side network takes advantage of the LPM by utilizing intermediate features from the frozen pre-trained model. As depicted in part (a) of Fig. 3, the lightweight ViT on the right is a scaled-down variant of the LPM on the left in terms of hidden dimensions and the number of layers, serving as the side network. The hidden dimensions of the side network are $1/k$ times that of the original, and its parameter weights are initialized randomly. Here, $k$ represents a reduction factor. The side network is shallower than the LPM in terms of the number of layers.
To utilize knowledge from the LPM, a novel fusion module $\mathcal{F}$ is introduced to fuse the intermediate features of the LPM and the side network at each layer, which will be elaborated upon in the subsequent section. Specifically, given the LPM $M$ with $L_{M}$ layers and the side network $S$ with $L_{S}$ layers, respectively parameterized by $\theta_{M}$ and $\theta_{S}$, the intermediate activation $\bm{z}_{S}^{l}$ of layer $l$ in the side network is obtained as follows:
@@ 3.2 Side Tuning
3.3 Asymmetric Input
Fine-grained information holds significant importance in the context of medical image analysis. To extract such detailed information, one straightforward approach involves increasing the resolution of the input image. However, due to the mechanics of input patchification and the self-attention layer in the ViT architecture, increasing the input resolution leads to a dramatic rise in memory consumption and training cost. For instance, in a ViT with a patch size of 16, increasing the input resolution from 224 to 512 results in an expansion of the input sequence length from 196 to 1024. Furthermore, the dimensions of the self-attention maps in the transformer block escalate from $196^{2}$ to $1024^{2}$. This considerable training overhead presents challenges for fine-tuning on high-resolution datasets.
Although high-resolution inputs make the training cost unfeasible, employing a pre-trained model for inference with high-resolution images remains practical. Consequently, we propose an asymmetric input strategy within the FPT+ framework. Specifically, the input images $I_{h}$ for the frozen pre-trained model are of high resolution, whereas the images fed to the learnable side network are low-resolution images $I_{l}$ that get down-sampled from the high-resolution ones. Therefore, the Equation 5 becomes
@@ 3.3 Asymmetric Input
3.4 Fine-grained Prompts and Fusion Modules
In this section, we introduce our proposed fusion modules $\mathcal{F}$ and fine-grained prompts. To fuse features from the LPM and the side network, a simple approach is to use an MLP layer to process and align the hidden dimension of these features with that of the side network, and then concatenate these features into the intermediate sequence of the side network. However, this approach will dramatically increase the length of the token sequence of the side network, significantly increasing memory consumption. Furthermore, popular LMPs are generally pre-trained on natural images, and since the pre-trained model remains frozen during training, integrating this out-of-domain knowledge might not seamlessly adapt to new medical downstream tasks. Moreover, directly fusing features across diverse scales and domains poses challenges for the side network’s ability to effectively learn from its input images.
To tackle these challenges, as illustrated in part (b) of Fig. 3, we introduce a small set of learnable embeddings $\bm{p}$ named fine-grained prompts and utilize a cross-attention mechanism (Chen et al., 2021; Vaswani et al., 2017) to summarize pre-trained knowledge from the LPM. Unlike prompt tuning (Lester et al., 2021; Jia et al., 2022) which directly uses prompts as part of the input sequence, our proposed fine-grained prompts serve as a bridge linking the frozen LPM and the side network. Specifically, in the context of cross-attention, we reuse the key $K_{M}$ and the value $V_{M}$ from the self-attention layer of the LPM, while the fine-grained prompts serve as the query. The cross-attention module allows the prompts to retrieve tokens of interest and summarize fine-grained information.
@@ 3.4 Fine-grained Prompts and Fusion Modules
where notations for the input source and the layer index are omitted for simplification. The $\text{CA}(\cdot)$ and $\text{LN}(\cdot)$ denote the cross-attention module and the layer normalization operation, respectively. The [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation, and the out-projector $f_{out}$ is a linear layer to align the hidden dimension of the fine-grained prompts to that of the side network. The out-projector is shared across all fusion modules to further increase the efficiency of FPT+. Notably, the number of the fine-grained prompts is considerably smaller than that of the input sequence, leading to reduced training cost. They are removed after each layer’s forward processing in the side network to maintain consistency of the sequence length across layers.
3.5 Important Tokens Selection
@@
In such context, we present a straightforward yet effective approach to mitigate the training expenses: we select important tokens, and only features associated with those important tokens are passed to the fusion modules. Specifically, we assess the tokens’ importance based on the self-attention maps (Dosovitskiy et al., 2021) in the LPM. The importance score of a token is defined as the average attention score of other tokens upon it. Subsequently, we select the top $m\%$ tokens with the highest scores as the important tokens, which are delivered to the side network through the fusion modules. This strategy alleviates the overhead introduced by the lengthy tokens from the high-resolution inputs, while preserving the discriminative features essential for downstream tasks.
3.6 Fine-grained Features Preloading
To further enhance the training efficiency, we choose not to employ any data augmentation for the high-resolution input of the frozen LPM, ensuring that features associated with an image of interest remain constant throughout training. Consequently, it becomes feasible to pre-store these features prior to training, resulting in notable reductions in the training cost. Leveraging the benefits of important tokens selection, the demand for storage and the overhead linked to data loading are also considerably reduced, thereby collectively accelerating the training process of FPT+. Although the high-resolution input of the LPM remains fixed, we apply data augmentation to the low-resolution input of the side network to maintain the diversity of the training samples. Note that the preloading technique is exclusive to FPT+, because the asymmetric input design allows for the application of preloading on the high-resolution input while avoiding overfitting and performance degradation through data augmentation on the low-resolution input.
Table 1: Comparison results with state-of-the-art PETL methods across all evaluation datasets. ’Params’ denotes the ratio of learnable parameters to the total number of parameters. ’Mem.’ denotes the memory usage (MB). ’Avg.’, ’PPE’, and ’PME’ denote the average performance across all datasets, performance-parameter-efficiency, and performance-memory-efficiency metrics. The best and second-best results are indicated in bold and underlined formatting, respectively.
@@ 3.6 Fine-grained Features Preloading
Method Params. Mem. DDR Messidor-2 ISIC2018 ISIC2016 DDSM CMMD COVID CHNCXR Avg. PPE PME
@@
Full fine-tuning 100 23128 89.79 ± 0.76 86.87 ± 0.53 96.65 ± 0.29 83.22 ± 1.02 92.49 ± 0.34 66.26 ± 0.69 99.85 ± 0.05 95.40 ± 1.05 88.82 65.73 65.73
Linear probing 0.01 3416 80.17 ± 0.85 79.73 ± 0.84 93.37 ± 0.31 81.27 ± 0.70 80.89 ± 0.39 60.23 ± 0.62 99.21 ± 0.07 91.34 ± 0.20 83.28 83.28 78.44
Prompt-tuning (Jia et al., 2022) 0.17 20582 81.07 ± 0.59 80.02 ± 2.34 94.20 ± 0.20 81.66 ± 1.03 82.67 ± 0.34 59.62 ± 0.83 99.27 ± 0.03 90.95 ± 0.82 83.68 83.62 63.47
Attention-tuning (Touvron et al., 2022) 33.8 20792 81.99 ± 1.85 81.87 ± 1.42 94.40 ± 0.40 80.11 ± 1.92 80.67 ± 2.33 60.03 ± 1.03 99.58 ± 0.19 92.07 ± 2.48 83.68 73.74 63.34
Adapter (Chen et al., 2022) 2.03 19360 85.84 ± 2.14 80.77 ± 7.48 95.96 ± 0.13 79.22 ± 2.26 80.76 ± 3.38 61.39 ± 1.40 99.17 ± 0.48 90.02 ± 0.91 84.14 83.41 64.61
BitFit (Ben-Zaken et al., 2022) 0.12 20382 84.31 ± 0.50 83.81 ± 1.11 94.84 ± 0.15 82.35 ± 0.62 84.77 ± 0.54 63.84 ± 1.80 99.81 ± 0.03 91.62 ± 0.95 85.67 85.63 65.11
LoRA (Hu et al., 2022) 0.68 20970 85.84 ± 2.42 86.08 ± 0.95 95.02 ± 0.22 84.22 ± 1.98 82.26 ± 4.04 62.26 ± 1.66 99.69 ± 0.05 92.13 ± 1.12 85.94 85.69 64.93
AdaLoRA (Zhang et al., 2023a) 0.52 20960 85.21 ± 0.34 86.07 ± 0.74 95.85 ± 0.24 84.23 ± 0.65 86.91 ± 0.30 62.45 ± 1.46 99.75 ± 0.04 92.25 ± 0.70 86.59 86.40 65.43
FPT (Huang et al., 2024a) 1.81 1824 84.13 ± 1.06 84.95 ± 2.01 93.88 ± 0.60 80.68 ± 1.13 90.52 ± 0.59 65.35 ± 0.94 99.70 ± 0.30 91.97 ± 1.36 86.40 85.73 83.60
@@
Full fine-tuning 100 23128 87.33 ± 0.77 84.06 ± 1.29 96.44 ± 0.26 84.46 ± 1.55 91.19 ± 0.32 64.70 ± 0.32 99.92 ± 0.03 95.31 ± 1.19 87.93 65.07 65.07
Linear probing 0.01 3416 80.77 ± 0.67 81.71 ± 0.69 93.44 ± 0.16 83.96 ± 0.93 81.06 ± 0.46 63.64 ± 1.08 99.40 ± 0.01 91.84 ± 0.18 84.48 84.48 79.57
Prompt-tuning (Jia et al., 2022) 0.17 20582 80.48 ± 0.34 80.71 ± 1.85 93.76 ± 0.11 83.41 ± 1.40 81.86 ± 0.34 63.30 ± 1.25 99.44 ± 0.14 91.97 ± 0.33 84.37 84.31 63.99
Attention-tuning (Touvron et al., 2022) 33.8 20792 82.38 ± 0.93 77.11 ± 3.29 92.60 ± 0.74 76.32 ± 4.05 75.58 ± 8.26 51.31 ± 2.63 99.57 ± 0.14 95.03 ± 0.30 81.24 71.59 61.49
Adapter (Chen et al., 2022) 2.03 19360 87.95 ± 0.54 84.99 ± 1.89 96.02 ± 0.13 82.30 ± 2.26 77.61 ± 8.86 57.56 ± 1.48 99.79 ± 0.12 89.73 ± 4.55 84.49 83.76 64.88
BitFit (Ben-Zaken et al., 2022) 0.12 20382 86.35 ± 0.42 84.68 ± 0.42 95.11 ± 0.15 84.60 ± 1.13 84.27 ± 0.34 66.03 ± 1.53 99.82 ± 0.04 93.43 ± 0.64 86.79 86.74 65.96
LoRA (Hu et al., 2022) 0.68 20970 85.14 ± 3.86 84.35 ± 1.27 95.97 ± 0.40 81.06 ± 6.82 $7$ 1.92 60.59 ± 2.36 99.89 ± 0.03 95.02 ± 0.70 84.97 84.72 64.20
AdaLoRA (Zhang et al., 2023a) 0.52 20960 86.39 ± 0.37 84.27 ± 0.66 95.79 ± 0.25 83.74 ± 0.88 88.16 ± 0.35 64.39 ± 0.14 99.64 ± 0.08 93.29 ± 0.79 86.96 86.76 65.71
FPT (Huang et al., 2024a) 1.81 1824 83.79 ± 0.76 84.69 ± 1.06 94.27 ± 0.18 83.05 ± 0.85 90.08 ± 0.68 66.23 ± 0.46 99.72 ± 0.05 91.55 ± 0.81 86.67 86.00 83.86
@@
4 Experiments
4.1 Datasets
We evaluate FPT+ using eight medical image datasets for classification tasks across four modalities, including fundus images (DDR (Li et al., 2019) and Messidor-2 (Decencière et al., 2014)), dermoscopic images (ISIC 2016 (Gutman et al., 2016) and ISIC 2018 (Codella et al., 2019)), mammography images (DDSM (Lee et al., 2017) and CMMD (Cui et al., 2021)), and chest X-ray images (COVID (Siddhartha, 2021) and CHNCXR (Jaeger et al., 2014)). The dataset sizes range from 662 to 12,522 samples, with classification categories varying from 2 to 7. When official splits for training, validation, and testing are available, they are adopted. Otherwise, a random partition of 70%/10%/20% is employed for training, validation, and testing, respectively.
4.2 Training and Evaluation Setup
4.2.1 Experiment Setup
In this study, we employ two popular variants of ViT (Dosovitskiy et al., 2021), specifically ViT-B(ase) and ViT-L(arge), as the large-scale pre-trained models, with 86 million and 304 million parameters, respectively. To evaluate the generalizability of different PETL methods, we use two pre-trained weights: one pre-trained in a supervised learning manner and the other using self-supervised learning with DINO (Caron et al., 2021). The pre-training dataset, ImageNet-21K (Ridnik et al., 2021), has been widely used for its robust generalizability with respect to natural images. The patch size of ViTs used in this work is 16. All methods are trained for 20 epochs with a mini-batch size of 16 on an NVIDIA RTX A6000 GPU (48GB) using PyTorch of version 2.0.0 (Paszke et al., 2017). We adopt a cosine annealing schedule for the learning rate during training. ViTs are trained using the AdamW optimizer (Loshchilov and Hutter, 2019) with cross-entropy as the loss function for all datasets. Each experiment is repeated 5 times, and we report the mean ± standard deviation values.

### v2

@@ III-A Vision Transformer
Within each self-attention layer, individual tokens in the input sequence are mapped to Q (query), K (key), and V (value) representations for computations in the attention module (Attn) [46] as follows:
@@
III-B Side Tuning
Unlike popular PETL methods which introduce supplementary modules into transformer blocks, our approach, drawing inspiration from [50, 44], adopts a lightweight and separate side network. The side network takes advantages of the LPM by utilizing intermediate features from the frozen pre-trained model. As depicted in part (a) of Fig. 3, the lightweight ViT on the right is a scaled-down variant of the LPM on the left in terms of hidden dimensions and the number of layers, serving as the side network. The hidden dimensions of the side network are $1/k$ times that of the original, and its parameter weights are initialized randomly. Here, $k$ represents a reduction factor. The side network is shallower than the LPM in terms of the number of layers.
To utilize knowledge from the LPM, a novel fusion module $\mathcal{F}$ is introduced to fuse the intermediate features of the LPM and those of the side network at each layer, which will be elaborated upon in the subsequent section. Specifically, given the LPM $M$ with $L_{M}$ layers and the side network $S$ with $L_{S}$ layers, respectively parameterized by $\theta_{M}$ and $\theta_{S}$, the intermediate activation $\bm{z}_{S}^{l}$ of layer $l$ in the side network is obtained as follows:
@@ III-B Side Tuning
III-C Asymmetric Input
Fine-grained information holds significant importance in medical image analysis. To extract such detailed information, one straightforward approach involves increasing the resolution of the input image. However, due to the mechanics of input patchification and the self-attention layer in the ViT architecture, increasing the input resolution leads to a dramatic rise in memory consumption and training cost. For instance, in a ViT with a patch size of 16, increasing the input resolution from 224 to 512 results in an expansion of the input sequence length from 196 to 1024. Furthermore, the dimensions of the self-attention maps in the transformer block escalate from $196^{2}$ to $1024^{2}$. This considerable training overhead presents challenges for fine-tuning on high-resolution datasets.
Although high-resolution inputs make the training cost unfeasible, employing a pre-trained model for inference with high-resolution images remains practical. We here propose an asymmetric input strategy within the FPT+ framework. Specifically, the input images $I_{h}$ for the frozen pre-trained model are of high resolution, whereas the images fed to the learnable side network are low-resolution images $I_{l}$ that get down-sampled from the high-resolution ones. Therefore, the Equation 5 becomes
@@ III-C Asymmetric Input
III-D Fine-grained Prompts and Fusion Modules
In this section, we introduce our proposed fusion modules $\mathcal{F}$ and fine-grained prompts. To fuse features from the LPM and those from the side network, a simple approach is to use an MLP layer to process and align the hidden dimensions of those features, and then concatenate them into the intermediate sequence of the side network. However, this approach will dramatically increase the length of the token sequence of the side network, significantly increasing memory consumption. Furthermore, popular LMPs are generally pre-trained on natural images, and since the pre-trained model remains frozen during training, integrating this out-of-domain knowledge might not seamlessly adapt to new medical downstream tasks. Moreover, directly fusing features across diverse scales and domains poses challenges for the side network’s ability to effectively learn from its input images.
To tackle these challenges, as illustrated in part (b) of Fig. 3, we introduce a small set of learnable embeddings $\bm{p}$ named fine-grained prompts and utilize a cross-attention mechanism [6, 46] to summarize pre-trained knowledge from the LPM. Unlike prompt tuning [31, 27] which directly uses prompts as part of the input sequence, our proposed fine-grained prompts serve as a bridge linking the frozen LPM and the side network. Specifically, in the context of cross-attention, we reuse the key $K_{M}$ and the value $V_{M}$ from the self-attention layer of the LPM, while the fine-grained prompts serve as the query. The cross-attention module allows the prompts to retrieve tokens of interest and summarize fine-grained information.
@@ III-D Fine-grained Prompts and Fusion Modules
where notations for the input source and the layer index are omitted for simplification. The $\text{CA}(\cdot)$ and $\text{LN}(\cdot)$ denote the cross-attention module and the layer normalization operation, respectively. The [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the concatenation operation, and the out-projector $f_{out}$ is a linear layer to align the hidden dimension of the fine-grained prompts to that of the side network. The out-projector is shared across all fusion modules to further increase efficiency. Notably, the number of the fine-grained prompts is considerably smaller than that of the input sequence, leading to reduced training cost. They are removed after each layer’s forward processing in the side network to maintain consistency of the sequence length across layers.
III-E Important Tokens Selection
@@
In such context, we present a straightforward yet effective approach to mitigate the training expenses: we select important tokens, and only features associated with those important tokens are passed to the fusion modules. Specifically, we assess the tokens’ importance based on the self-attention maps [15] in the LPM. The importance score of a token is defined as the average attention score of other tokens upon it. Subsequently, we select the top $m\%$ tokens with the highest scores as the important tokens, which are delivered to the side network through the fusion modules. This strategy alleviates the overhead introduced by the lengthy tokens from the high-resolution inputs, while preserving the discriminative features essential for downstream tasks.
III-F Fine-grained Features Preloading
To further enhance the training efficiency, we choose not to employ any data augmentation for the high-resolution input of the frozen LPM, ensuring that features associated with an image of interest remain constant throughout training. Consequently, it becomes feasible to pre-store these features prior to training, resulting in notable reductions in the training cost. Leveraging the benefits of important tokens selection, the demand for storage and the overhead linked to data loading are also considerably reduced, thereby collectively accelerating the training process. Although the high-resolution input of the LPM remains fixed, we apply data augmentation to the low-resolution input of the side network to maintain the diversity of the training samples. Note that the preloading technique is exclusive to FPT+, because the asymmetric input design allows for the application of preloading on the high-resolution input while avoiding overfitting and performance degradation through data augmentation on the low-resolution input.
TABLE I: Comparison results with state-of-the-art PETL methods across all evaluation datasets. ’Params’ denotes the ratio of learnable parameters to the total number of parameters. ’Mem.’ denotes the memory usage (MB). ’Avg.’, ’PPE’, and ’PME’ denote the average performance across all datasets, performance-parameter-efficiency, and performance-memory-efficiency metrics. The best and second-best results are indicated in bold and underlined formatting, respectively.
@@ III-F Fine-grained Features Preloading
Method Params. $\downarrow$ Mem. $\downarrow$ DDR Messidor-2 ISIC2018 ISIC2016 DDSM CMMD COVID CHNCXR Avg. $\uparrow$ PPE $\uparrow$ PME $\uparrow$
@@
Full fine-tuning 100 23,128 89.79 ± 0.76 86.87 ± 0.53 96.65 ± 0.29 83.22 ± 1.02 92.49 ± 0.34 66.26 ± 0.69 99.85 ± 0.05 95.40 ± 1.05 88.82 65.73 65.73
Linear probing 0.01 3,416 80.17 ± 0.85 79.73 ± 0.84 93.37 ± 0.31 81.27 ± 0.70 80.89 ± 0.39 60.23 ± 0.62 99.21 ± 0.07 91.34 ± 0.20 83.28 83.28 78.44
Prompt-tuning [27] 0.17 20,582 81.07 ± 0.59 80.02 ± 2.34 94.20 ± 0.20 81.66 ± 1.03 82.67 ± 0.34 59.62 ± 0.83 99.27 ± 0.03 90.95 ± 0.82 83.68 83.62 63.47
Attention-tuning [45] 33.8 20,792 81.99 ± 1.85 81.87 ± 1.42 94.40 ± 0.40 80.11 ± 1.92 80.67 ± 2.33 60.03 ± 1.03 99.58 ± 0.19 92.07 ± 2.48 83.68 73.74 63.34
Adapter [7] 2.03 19,360 85.84 ± 2.14 80.77 ± 7.48 95.96 ± 0.13 79.22 ± 2.26 80.76 ± 3.38 61.39 ± 1.40 99.17 ± 0.48 90.02 ± 0.91 84.14 83.41 64.61
BitFit [2] 0.12 20,382 84.31 ± 0.50 83.81 ± 1.11 94.84 ± 0.15 82.35 ± 0.62 84.77 ± 0.54 63.84 ± 1.80 99.81 ± 0.03 91.62 ± 0.95 85.67 85.63 65.11
LoRA [22] 0.68 20,970 85.84 ± 2.42 86.08 ± 0.95 95.02 ± 0.22 84.22 ± 1.98 82.26 ± 4.04 62.26 ± 1.66 99.69 ± 0.05 92.13 ± 1.12 85.94 85.69 64.93
AdaLoRA [51] 0.52 20,960 85.21 ± 0.34 86.07 ± 0.74 95.85 ± 0.24 84.23 ± 0.65 86.91 ± 0.30 62.45 ± 1.46 99.75 ± 0.04 92.25 ± 0.70 86.59 86.40 65.43
FPT [23] 1.81 1,824 84.13 ± 1.06 84.95 ± 2.01 93.88 ± 0.60 80.68 ± 1.13 90.52 ± 0.59 65.35 ± 0.94 99.70 ± 0.30 91.97 ± 1.36 86.40 85.73 83.60
@@
Full fine-tuning 100 23,128 87.33 ± 0.77 84.06 ± 1.29 96.44 ± 0.26 84.46 ± 1.55 91.19 ± 0.32 64.70 ± 0.32 99.92 ± 0.03 95.31 ± 1.19 87.93 65.07 65.07
Linear probing 0.01 3,416 80.77 ± 0.67 81.71 ± 0.69 93.44 ± 0.16 83.96 ± 0.93 81.06 ± 0.46 63.64 ± 1.08 99.40 ± 0.01 91.84 ± 0.18 84.48 84.48 79.57
Prompt-tuning [27] 0.17 20,582 80.48 ± 0.34 80.71 ± 1.85 93.76 ± 0.11 83.41 ± 1.40 81.86 ± 0.34 63.30 ± 1.25 99.44 ± 0.14 91.97 ± 0.33 84.37 84.31 63.99
Attention-tuning [45] 33.8 20,792 82.38 ± 0.93 77.11 ± 3.29 92.60 ± 0.74 76.32 ± 4.05 75.58 ± 8.26 51.31 ± 2.63 99.57 ± 0.14 95.03 ± 0.30 81.24 71.59 61.49
Adapter [7] 2.03 19,360 87.95 ± 0.54 84.99 ± 1.89 96.02 ± 0.13 82.30 ± 2.26 77.61 ± 8.86 57.56 ± 1.48 99.79 ± 0.12 89.73 ± 4.55 84.49 83.76 64.88
BitFit [2] 0.12 20,382 86.35 ± 0.42 84.68 ± 0.42 95.11 ± 0.15 84.60 ± 1.13 84.27 ± 0.34 66.03 ± 1.53 99.82 ± 0.04 93.43 ± 0.64 86.79 86.74 65.96
LoRA [22] 0.68 20,970 85.14 ± 3.86 84.35 ± 1.27 95.97 ± 0.40 81.06 ± 6.82 $7$ 1.92 60.59 ± 2.36 99.89 ± 0.03 95.02 ± 0.70 84.97 84.72 64.20
AdaLoRA [51] 0.52 20,960 86.39 ± 0.37 84.27 ± 0.66 95.79 ± 0.25 83.74 ± 0.88 88.16 ± 0.35 64.39 ± 0.14 99.64 ± 0.08 93.29 ± 0.79 86.96 86.76 65.71
FPT [23] 1.81 1, 824 83.79 ± 0.76 84.69 ± 1.06 94.27 ± 0.18 83.05 ± 0.85 90.08 ± 0.68 66.23 ± 0.46 99.72 ± 0.05 91.55 ± 0.81 86.67 86.00 83.86
@@
IV Experiments
IV-A Datasets
We evaluate FPT+ using eight medical image datasets for classification tasks across four modalities, including fundus images (DDR [33] and Messidor-2 [11]), dermoscopic images (ISIC 2016 [18] and ISIC 2018 [9]), mammography images (DDSM [30] and CMMD [10]), and chest X-ray images (COVID [42] and CHNCXR [26]). The dataset sizes range from 662 to 12,522 samples, with classification categories varying from 2 to 7. When official splits for training, validation, and testing are available, they are adopted. Otherwise, a random partition of 70%/10%/20% is employed for training, validation, and testing, respectively.
IV-B Training and Evaluation Setup
IV-B1 Experiment Setup
In this study, we employ two popular variants of ViT [15], specifically ViT-B(ase) and ViT-L(arge), as our large-scale pre-trained models, with 86 million and 304 million parameters, respectively. To evaluate the generalizability of different PETL methods, we experiment with two types of pre-trained weights: one obtained through supervised learning and the other via self-supervised learning using DINO [5]. These models, which are pre-trained on the ImageNet-21K dataset [40], are widely recognized for their robust generalizability in natural image tasks. The patch size of ViTs used in this work is 16. All methods are trained for 20 epochs with a mini-batch size of 16 on an NVIDIA RTX A6000 GPU (48GB) using PyTorch of version 2.0.0 [36]. We adopt a cosine annealing schedule for the learning rate during training. ViTs are trained using the AdamW optimizer with cross-entropy as the loss function for all datasets. Each experiment is repeated 5 times, and we report the mean ± standard deviation values.